IRR AND NPV A company is analyzing two mutually exclusive projects, S and L, with the following cash flows: Project s $1,000 $886.43 $260 $5 $5 Project L $1,000 $10 $240 $400 $819.73 The companys WACC is 9.0% what is the IRR of the better project? Hint: The better project may or may not be the one th e hi her 1 R und your answer to tvo decimal places
Answer #1

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

S:

Present value of inflows=886.43/1.09+260/1.09^2+5/1.09^3+5/1.09^4

=$1039.48

NPV=Present value of inflows-Present value of outflows

=$1039.48-$1000

=$39.48

L:

Present value of inflows=10/1.09+240/1.09^2+400/1.09^3+819.73/1.09^4

=$1100.77

NPV=Present value of inflows-Present value of outflows

=$1100.77-$1000

=$100.77

Hence L is a better project having higher NPV.

Let irr be x%
At irr,present value of inflows=present value of outflows.

1000=10/1.0x+240/1.0x^2+400/1.0x^3+819.73/1.0x^4

hence x=irr=12.20%(Approx).